# Towards Fair In-Context Learning with Tabular Foundation Models

### v3

@@ Appendix A Datasets
We experiment on tasks from the recently proposed folktables (Ding et al., 2022), which contains data extracted from the American Community Survey (ACS) Public Use Microdata Sample (PUMS) (Ding et al., 2022). More specifically, we experiment with the following ACS PUMS tasks:
@@
These tasks were selected to reflect a range of real-world predictive challenges with fairness concerns. We use the data of the year 2018 from the state of Alabama (AL), which is one of the states with the largest fairness violation (Ding et al., 2022)222We also perform experiments on data from other states, and observed that the results presented in the paper remain consistent..
A limitation of the ACS PUMS datasets is that they are US-centric; we diversify the experimental setup by including other tasks and datasets. Specifically, we also experiment on the following tabular datasets and tasks:
Diabetes (Gardner et al., 2023): The diabetes prediction task uses features related to physical health, lifestyle factors, and chronic conditions, derived from the BRFSS questionnaires. Demographic attributes like race, sex, state, and income are also included. The target is a binary indicator of whether the respondent has ever been diagnosed with diabetes.
German Credit (Frank, 2010): The German Credit dataset contains 20 attributes of 1,000 individuals. We create the task of classifying people according to whether they have a good or bad credit risk using age (over or below 25 years old) as the sensitive attribute.
CelebA (Liu et al., 2018): The dataset contains 202,599 samples described with 40 facial attributes of human annotated images. We create the task of predicting attractiveness with facial attributes using gender as the sensitive attribute (Kenfack et al., 2024b). Note that we do not train the model with images and consider this task to diversify the experimental tasks.
@@ B.1 Fairness Metrics
Demographic parity (DP): DP enforces equal positive outcome rate for different groups (Dwork et al., 2012) and is defined as follows:
@@
Equalized Odds (EOD): EOdds is satisfied when the model makes correct and incorrect predictions at the same rate for different demographic groups (Hardt et al., 2016). The metric enforces equal true positive and false positive rates across groups and is measured as follows;
@@ B.2 Correlation Remover
The Correlation Remover (Feldman et al., 2015) is a preprocessing technique designed to eliminate linear correlations between sensitive attributes and non-sensitive features in a dataset. This method is particularly useful in mitigating biases that may arise due to such correlations, especially when employing linear models.
@@ B.3 Uncertainty measure with conformal prediction
While any uncertainty measurement method could be used in our method, we employ conformal prediction due to its strong coverage guarantees. Specifically, we used Split Conformal Prediction (Angelopoulos et al., 2023), which is a distribution-free and model-agnostic method that provides prediction sets for classification tasks, ensuring a user-specified coverage level with no assumptions beyond data exchangeability.
@@
Consider a simplified task consisting of two non-sensitive ( and ) features one binary target () and a binary sensitive attribute ().
We first train a sensitive attribute classifier, using a fraction of the dataset (20% in our experiment), to estimate the uncertainties. We train the sensitive attribute classifier using and to predict .
@@
A conformal predictor (details provided above) will return for a given (test) sample a prediction set that contains the true sensitive attribute with probability of at least . For example, setting means 95% of the data will contain the true label in their prediction sets. The size of the prediction set, therefore, provides information about the prediction uncertainty, i.e., means the prediction is confident with at least probability, and means the prediction is uncertain. The intuition of our method is that using samples with uncertain predictions makes it challenging for the foundation model to infer and rely on the sensitive attributes to make predictions on the target.
@@
We therefore filter out a demonstration example if and select it if . helps to control the fairness-accuracy tradeoff since smaller corresponds to dropping more demonstration examples with higher confident sensitive attributes prediction.
Appendix C Supplementary resutls
| Dataset | ICL Method | TabPFN | TabICL | ||
| Accuracy | F1 Score | Accuracy | F1 Score | ||
| ACSIncome | Vanilla | ||||
| Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| ACSTravelTime | Vanilla | ||||
| Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| ACSPublicCoverage | Vanilla | ||||
| Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| ACSEmployment | Vanilla | ||||
| Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| ACSMobility | Vanilla | ||||
| Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| German Credit | Vanilla | ||||
| Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| Diabetes | Vanilla | ||||
| Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| CelebA | Vanilla | ||||
| Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
The results in Figure 1 and Table 5 show that improving fairness in the ACSPublicCoverage and Diabetes datasets is particularly challenging. We observe that unfairness is already pretty small, and the fairness methods alter the training data, either by transforming the features (correlation removal) or reducing its size (group-balanced and Uncertain methods), which reduces accuracy, without necessarily improving on fairness. For Table 1, we used a fixed value for the hyperparameter controlling the fairness accuracy tradeoff, for the uncertain method and for correlation removal. Figure 7(c) and 7(d) show that for different values of , there are Pareto-optimality points of the uncertain method that significantly reduce unfairness compared to random demonstration selection baselines. These points correspond to values of . In sum, for some tasks, in a low unfairness regime such as PublicCoverage and Diabetes, smaller values of are needed to reduce unfairness, which means more data points with a confident prediction of the sensitive attribute must be removed from the demonstration set. The value used in practice can be defined by stakeholders, decision-makers or business needs, considering that smaller means increased fairness and a drop in accuracy, and higher means a decrease in fairness and higher accuracy.
| Dataset | ICL Method | Accuracy | DP | EOP | EOD |
| ACSIncome | Vanilla | ||||
| Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| ACSEmployment | Vanilla | ||||
| Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| ACSPublicCoverage | Vanilla | ||||
| Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| ACSTravelTime | Vanilla | ||||
| Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| ACSMobility | Vanilla | ||||
| Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| Diabetes | Vanilla | ||||
| Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| German Credit | Vanilla | ||||
| Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| CelebA | Vanilla | ||||
| Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN |

### v4

@@ Appendix A Datasets
We experiment on tasks from the recently proposed folktables (ding_retiring_2022), which contains data extracted from the American Community Survey (ACS) Public Use Microdata Sample (PUMS) (ding_retiring_2022). More specifically, we experiment with the following ACS PUMS tasks:
@@
These tasks were selected to reflect a range of real-world predictive challenges with fairness concerns. A limitation of the ACS PUMS datasets is that they are US-centric; we diversify the experimental setup by including other tasks and datasets. Specifically, we also evaluate on the following tabular datasets and tasks:
Diabetes (gardner2023tableshift): The diabetes prediction task uses features related to physical health, lifestyle factors, and chronic conditions, derived from the BRFSS questionnaires. Demographic attributes like race, sex, state, and income are also included. The target is a binary indicator of whether the respondent has ever been diagnosed with diabetes.
German Credit (frank2010uci): The German Credit dataset contains 20 attributes of 1,000 individuals. We create the task of classifying people according to whether they have a good or bad credit risk using age (over or below 25 years old) as the sensitive attribute.
CelebA (liu2018large): The dataset contains 202,599 samples described with 40 facial attributes of human annotated images. We create the task of predicting attractiveness with facial attributes using gender as the sensitive attribute (kenfack2024fairness). Note that we do not train the model with images and consider this task to diversify the experimental tasks.
@@ B.1 Fairness Metrics
Demographic parity (DP): DP enforces equal positive outcome rate for different groups (dwork_fairness_2012) and is defined as follows:
@@
Equalized Odds (EOD): EOdds is satisfied when the model makes correct and incorrect predictions at the same rate for different demographic groups (hardt_equality_2016). The metric enforces equal true positive and false positive rates across groups and is measured as follows;
@@ B.2 Correlation Remover
The Correlation Remover (feldman2015certifying) is a preprocessing technique designed to eliminate linear correlations between sensitive attributes and non-sensitive features in a dataset. This method is particularly useful in mitigating biases that may arise due to such correlations, especially when employing linear models.
@@ B.3 Uncertainty measure with conformal prediction
While any uncertainty measurement method could be used in our method, we employ conformal prediction due to its strong coverage guarantees. Specifically, we used Split Conformal Prediction (angelopoulos2023conformal), which is a distribution-free and model-agnostic method that provides prediction sets for classification tasks, ensuring a user-specified coverage level with no assumptions beyond data exchangeability.
@@
Consider a simplified task consisting of two non-sensitive features ( and ), one binary target (), and a binary sensitive attribute ().
@@
To estimate the uncertainties, we first train a sensitive attribute classifier, using a fraction of the dataset (20% in our experiments). We train the sensitive attribute classifier (Logistic Regression or TabPFN) using and to predict . A conformal predictor returns for a given (test) sample a prediction set that contains the true sensitive attribute with probability of at least . For example, setting means 95% of the data will contain the true label in their prediction sets. The size of the prediction set, therefore, provides information about the prediction uncertainty, i.e., means the prediction is confident with at least probability, and means the prediction is uncertain. The intuition of our method is that using samples with uncertain predictions makes it challenging for the foundation model to infer and rely on the sensitive attributes to make predictions on the target.
@@
We therefore include a training example in the context set if and filter it out otherwise. helps to control the fairness-utility tradeoff since a smaller corresponds to filtering out more demonstration examples with higher confidence sensitive attributes prediction, which can better improve fairness while impacting accuracy as the size of the in-context set is reduced.
Appendix C Runtime Comparison of Fairness Interventions
In addition to fairness–utility tradeoffs, we provide a detailed comparison of the computational cost of the fairness interventions across tabular foundation models. While prior sections focused on predictive performance, practical deployment also requires understanding efficiency, particularly in large-scale or resource-constrained settings.
Setup.
We measured wall-clock running time on an NVIDIA A100 GPU (20 GB memory). Each run includes preprocessing, fairness intervention, and in-context learning inference using the full training set and test set. Results are averaged over three seeds, with standard deviations reported. The largest dataset in our experiments, CelebA, was used to provide representative estimates of runtime under high load. Table 5 summarizes runtimes across methods and models. We make the following observations: (i) TabPFN is the most efficient (8–47s), benefiting from its compact architecture. The uncertainty-based variant with LR is fastest, while using TabPFN itself for uncertainty estimation adds moderate overhead. (ii) TabICL shows moderate cost (57–97s), with relatively small variance across interventions. (iii) TabDPT is substantially slower ( 595s), reflecting the higher computational demands of its transformer-based architecture.
Across models, uncertainty-based with Logistic Regression is consistently the fastest intervention, while correlation removal and group balancing introduce negligible overhead compared to vanilla inference. Recall TabPFN uses at most 10K samples, which reduces the context size compared to other models, thereby explaining the faster inference. Likewise, the uncertainty-based intervention reduces the context size compared to other interventions. Running time on the ACSIncome dataset (Table 6) confirms that fairness interventions themselves are computationally inexpensive; the runtime is dominated by the choice of tabular foundation model. For practitioners, TabPFN with uncertainty-based sampling (LR) offers the best balance of fairness gains and efficiency, while TabDPT provides stronger model capacity at substantially higher cost.
| Model | Vanilla | Group Balanced | Correlation Remover | Uncertain+LR | Uncertain+TabPFN |
| TabDPT | 596.04 69.12 | 594.98 68.35 | 596.83 70.10 | 595.58 65.73 | 316.86 50.17 |
| TabICL | 69.50 0.75 | 67.84 0.85 | 70.11 1.26 | 57.05 1.24 | 96.65 1.06 |
| TabPFN | 10.71 1.27 | 10.84 1.16 | 13.94 1.35 | 8.45 0.82 | 47.18 0.65 |
| Model | Vanilla | Group Balanced | Correlation Remover | Uncertain+LR | Uncertain+TabPFN |
| TabDPT | 154.35 53.13 | 153.23 53.06 | 151.75 52.88 | 155.09 53.41 | 303.63 52.21 |
| TabICL | 45.05 0.62 | 45.23 0.04 | 44.63 0.62 | 45.09 0.29 | 45.43 1.40 |
| TabPFN | 1.85 0.87 | 1.96 1.07 | 3.22 0.32 | 1.87 0.91 | 3.65 0.76 |
Appendix D Supplementary resutls